The following information concerns production in the Baking Department for December. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Dec. 1 Bal., 5,700 units, 1/3 completed 15,200
31 Direct materials, 102,600 units 205,200 220,400
31 Direct labor 60,320 280,720
31 Factory overhead 33,928 314,648
31 Goods finished, 104,100 units 303,980 10,668
31 Bal., ? units, 3/5 completed 10,668

a. Determine the number of units in work in process inventory at December 31.
units

b. Determine the equivalent units of production for direct materials and conversion costs in December. If an amount is zero, enter in "0".

Baking Department
Equivalent Units of Production for Direct Materials and Conversion Costs
For December
Whole Units Equivalent Units
Direct Materials
Equivalent
Units Conversion
Inventory in process, December 1
Started and completed in December
Transferred to finished goods in December
Inventory in process, December 31
Total
Answer #1
a
Number of units in work in process inventory at December 31 4200 units =5700+102600-104100
b
Baking Department
Equivalent Units of Production for Direct Materials and Conversion Costs
For December
Whole Units Equivalent Units Equivalent
Direct Materials Units Conversion
Inventory in process, December 1 5700 0 3800 =5700*2/3
Started and completed in December 98400 98400 98400 =104100-5700
Transferred to finished goods in December 104100 98400 102200
Inventory in process, December 31 4200 4200 2520 =4200*3/5
Total 108300 102600 104720